# Data-Model Co-Evolution: Growing Test Sets to Refine LLM Behavior

### v1

@@ Abstract.
A long-standing challenge in machine learning has been the rigid separation between data work and model refinement, enforced by slow fine-tuning cycles. The rise of Large Language Models (LLMs) overcomes this historical barrier, allowing applications developers to instantly govern model behavior by editing prompt instructions. This shift enables a new paradigm: data-model co-evolution, where a living test set and a model’s instructions evolve in tandem. We operationalize this paradigm in an interactive system designed to address the critical challenge of encoding subtle, domain-specific policies into prompt instructions. The system’s structured workflow guides people to discover edge cases, articulate rationales for desired behavior, and iteratively evaluate instruction revisions against a growing test set. A user study shows our workflow helps participants refine instructions systematically and specify ambiguous policies more concretely. This work points toward more robust and responsible LLM applications through human-in-the-loop development aligned with local preferences and policies.
@@
$c$s: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language generation
Figure 1. Overview of the data-model co-evolution paradigm. Traditionally, prompt instructions (model) and test set examples (data) have been developed separately with slower model tuning cycles. With rapid iteration, however, these two artifacts can now evolve together. Our approach operationalizes this opportunity by treating each discovered failure as a test case and validating each prompt instruction revision against the growing test set, enabling systematic refinement.
@@ 1. Introduction
Large Language Models (LLMs) are driving a new wave of application development. This empowers people, including those without programming expertise, to direct model behavior through natural language instructions, a practice known as prompt engineering (Zamfirescu-Pereira et al., 2023; Subramonyam et al., 2025). However, it comes with a critical challenge: a general-purpose model inherently lacks the local context and nuanced policies of a specific application (Palla et al., 2025; In et al., 2025). For instance, a small online community may wish to foster playful banter by allowing specific slang, a nuance lost on a strict, one-size-fits-all content moderation filter from a major tech vendor, such as OpenAI and Google. This gap forces people into a reactive cycle of prompt tinkering, making small, ad-hoc edits and spot-checking them against a few examples to patch failures at the margins.
@@
This ad-hoc approach reflects a deeper mismatch between the new speed of development and the old tools at hand. Many current practices for evaluation, such as building static test sets, are holdovers from the traditional machine learning era (Sambasivan et al., 2021). Back then, the slow, costly process of model fine-tuning necessitated a strict separation between data curation and model refinement. In today’s interactive landscape, where application developers can iterate on prompts in seconds, this inherited separation has become a major source of friction, preventing a truly fluid and responsive development process.
@@
This friction is compounded by a new bottleneck: the sheer difficulty of authoring comprehensive test data for the long tail of modern LLM-based applications. Manually anticipating the diverse ways users will interact with generative models is a formidable task. To compensate, specialized and labor-intensive efforts have emerged. Red teaming, for example, uncovers critical vulnerabilities but often operates as a siloed, episodic exercise whose findings rarely integrate back into a persistent test suite (Feffer et al., 2024; Kiela et al., 2021; Quaye et al., 2024). Similarly, policy specifications or “constitutions” provide high-level guidance but tend to be static artifacts, difficult to update against a continuous stream of new edge cases (Bai et al., 2022; Petridis et al., 2024). Whether through ad-hoc tinkering, episodic testing, or static policies, a crucial feedback loop between discovering failures and refining behavior remains broken. We argue that the problem is the separation of data and model, where prompting is fast, but evaluation is slow.
To address this, we introduce data–model co-evolution, a new paradigm where the instructions (the model) and the test examples (the data) are treated as intertwined artifacts that grow together. Crucially, we define the ‘model’ from the application developer’s perspective: it is the editable specification, typically a system prompt that encodes policies and instructions that the part of the system they can actually edit to control its behavior. In our approach, this model helps discover its own blind spots, which are then immediately transformed into persistent test cases. Each revision is then tested against this growing set of real-world examples, creating the tight feedback loop that current practices lack.
@@
To operationalize this paradigm, we present an interactive system that supports users to refine prompt instructions and create test sets together for generative applications. Users begin with an initial policy and prompt. The system proposes candidate user inputs, including those likely to be ambiguous under the given instruction, and runs the target model to reveal behavior. When the output is incorrect, the system engages the user in a dialog: they mark the example and articulate a short rationale. To test generalization to other examples, the system generates small neighborhoods of similar cases for quick labeling. These labeled slices both grow the test set and inform revisions to the prompt. Users are then directed to compare model responses before and after revision, with an LLM-as-judge (Zheng et al., 2023) seeded by the user’s own labels. Crucially, all labeled data persist, yielding a living, incrementally expanding test set.
@@
Our approach is grounded in formative interviews with five practitioners, which revealed a clear need for workflows that better integrate systematic, data-driven evaluation into the development loop. We then evaluated our system in a controlled study (N=12) against a baseline. Our workflow led participants to produce more detailed instructions with explicit boundaries and examples, build larger and richer test sets rapidly and systematically, induced more consistent changes in model behavior, and reported higher satisfaction compared to a baseline.
@@
  • A new paradigm of data–model co-evolution, positioning test-set growth and specification refinement as incrementally co-evolving artifacts in LLM application development, informed by formative interviews with practitioners.
  • An interactive system that embodies this paradigm through structured workflows that support edge case discovery, rationale elicitation, neighborhood probing, and regression testing against a living test set.
  • Empirical insights from a controlled lab study (N=12), demonstrating that our workflow produces more specific prompt instructions, larger test sets, more consistent changes to model behavior, and higher user satisfaction than a baseline.
Our work contributes to HCI and human-AI interaction by advancing methods for democratizing LLM development and enabling more responsible, policy-aware use of AI. By shifting from prompt tinkering or episodic red teaming toward continuous, human-in-the-loop test-set growth, we show how interactive systems can help diverse developers—not only AI experts—make sense of subtle application-specific norms, codify them incrementally, and build more robust and accountable LLM applications.
@@ 2.1. Human-Centered Prompt and Policy Design
Recent HCI research on LLMs and prompt engineering has framed prompts as design artifacts that encode requirements and governance choices. For example, ConstitutionMaker (Petridis et al., 2024) demonstrates how user critiques can be transformed into enforceable principles, while ROPE (Ma et al., 2025) highlights the role of prompts as requirement specifications. Policy Maps (Lam et al., 2024) treats policy design as a cartographic practice, providing interactive maps that help practitioners navigate and demarcate regions of acceptable and unacceptable LLM behaviors. PromptHive (Reza et al., 2025) illustrates how domain experts can co-create prompts. Together, these projects emphasize that prompts are sites where human values and policies are concretized.
Complementing these policy-oriented perspectives, other work has built interactive systems to help practitioners explore, compare, and refine prompts. PromptIDE (Strobelt et al., 2022) provides visual analytics support for prompt iteration, EvalLM (Kim et al., 2024) introduces workflows for comparing different prompt versions through structured evaluation, and ChainForge (Arawjo et al., 2024) enables rapid experimentation with LLM prompt chains. In addition, researchers have explored alternative familiar form factors, spreadsheets, for supporting prompt iteration in the context of data labeling and classification (He et al., 2025; Qian et al., 2025).
In parallel, the Natural Language Processing (NLP) community has largely pursued automatic prompt optimization, treating prompt design as a black-box search problem (Cheng et al., 2024). Recent techniques include black-box optimization for alignment, multi-branch optimization strategies (Yang et al., 2024), and embedding-based prompt search (Cheng et al., 2023). These methods improve performance but rarely help stakeholders articulate policies or test boundary cases.
@@ 2.2. Interactive Machine Learning and Labeling
Interactive machine learning (IML) has long shown that humans and models can co-adapt through iterative data curation (Amershi et al., 2014; Dudley and Kristensson, 2018). The Machine Teaching paradigm argued for exposing teaching levers, so non-experts can shape model behavior with principled, design-forward workflows. (Simard et al., 2017).
Recent mixed-initiative tools leverage LLMs for data annotation. PaTAT (Gebreegziabher, 2025) supports qualitative coding by synthesizing interpretable rules from a user’s codes, and Wang et al. (Wang et al., 2024) proposes a collaboration workflow where LLMs produce labels and explanations, a verifier assesses their quality, and human annotators re-annotate lower-quality ones.
@@
In parallel, the data-centric AI movement argues that systematic data design and curation beyond model tuning drive performance, advocating practices for engineering data quality (Whang et al., 2023). Weak-supervision frameworks like Snorkel (Ratner et al., 2017) excel at throughput and coverage, but often miss opportunities for interactive elicitation.
@@ 2.3. Data Iteration and Test-Driven Development
A large body of HCI work argues that the importance of datasets in ML development. Chameleon (Hohman et al., 2020) shows how practitioners can attribute performance changes to data, comparing features, splits, and successive dataset versions through interactive visualizations. A series of interview studies further document the difficulty and iterative nature of data work in machine learning development in practice (Sambasivan et al., 2021; Holstein et al., 2019; Crisan and Fiore-Gartland, 2021; Qian et al., 2025).
Behavioral testing frameworks operationalize this stance by turning errors into executable tests. CheckList (Ribeiro et al., 2020) provides a tooling to ideate perturbations and generate broad suites of checks, showing that teams uncover substantially more actionable bugs than with ad-hoc evaluation. There also exist the use of a domain-specific language to define precise error groups and run counterfactual rewrites (Wu et al., 2019, 2020). Zeno (Cabrera et al., 2023a) develops an interactive web UI to slice data, specify behaviors, and track results across tasks. AdaTest (Ribeiro and Lundberg, 2022) takes human feedback to generate unit tests for debugging NLP models. We adopt the behavioral stance by creating slices from incorrectly generated examples, but center policy-boundary ambiguity with LLM generation, and let users compare new versions to older ones for assessment as well.
@@ 2.4. Bidirectional Human-AI Alignment and Criteria Drift
2.5. Human’s Roles in Red Teaming
Red teaming has emerged as a central strategy for stress-testing generative models. Automated pipelines exist for generating adversarial inputs (Perez et al., 2022), but researchers emphasize their limited creativity and coverage, making human input indispensable (Feffer et al., 2024). To scaffold human creativity at scale, the Adversarial Nibbler challenge (Quaye et al., 2024) recruited diverse communities to expose vulnerabilities in text-to-image generation, based on the Dynabench (Kiela et al., 2021) project. In addition, Yeh et al. (Yeh et al., 2025) and Deng et al. (Deng et al., 2025) leverages visual analytic solutions to help red-teaming professionals explore the space of adversarial prompts.
@@ 3. Formative Study
To ground our work in real-world practices, we conducted a formative study to understand the challenges industry practitioners face when developing and refining LLM-based applications. We sought to answer: how do teams currently discover, diagnose, and address failures in their LLM systems, and how do these failures feed into their processes for improving their applications?
@@ 3.1. Participants and Procedure
We conducted 60-minute semi-structured interviews over video conferencing with five practitioners from various industries. The participants came from diverse professional roles, including data scientists, software engineers, product managers, and AI researchers with two to eleven years of experience in their fields. They were actively building and deploying a range of LLM applications, including natural-language-to-SQL query systems, call center dialogue summarization and classification, banking chatbots, and AI-assisted planning applications. Our interviews focused on their day-to-day workflows for development, evaluation, and iteration. The interviews were recorded, transcribed, and coded through a thematic analysis. The study was approved by our institution’s IRB.
@@ 3.2.1. Development is a Process of Constant, Unending Iteration
This cycle of iteration–collecting new cases, updating prompts, and testing again–was described as never-ending. Below, we take a closer look at how practitioners actually conduct this iterative process.

### v2

# Data-Prompt Co-Evolution: Growing Test Sets to Refine LLM Behavior

@@ Abstract.
Large Language Models (LLMs) are increasingly embedded in applications, and people can shape model behavior by editing prompt instructions. Yet encoding subtle, domain-specific policies into prompts is challenging. Although this process often benefits from concrete test cases, test data and prompt instructions are typically developed as separate artifacts, reflecting traditional machine learning practices in which model tuning was slow and test sets were static. We argue that the fast, iterative nature of prompt engineering calls for removing this separation and enabling a new workflow: data-prompt co-evolution, where a living test set and prompt instructions evolve in tandem. We present an interactive system that operationalizes this workflow. It guides application developers to discover edge cases, articulate rationales for desired behavior, and iteratively evaluate revised prompts against a growing test set. A user study shows our workflow helps people refine prompts systematically, better aligning them with their intended policies. This work points toward more robust and responsible LLM applications through human-in-the-loop development.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791222isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language generation
Figure 1. Overview of the data-prompt co-evolution workflow. In common workflows for developing LLM-based applications, test data is typically treated as fixed and developed separately from prompt instructions. We propose a new interactive workflow in which test data and prompt instructions evolve together. Our approach helps users characterize failures, incorporate them into a growing test set that expands over iterations, and evaluate prompt revisions against this evolving set, enabling systematic, evidence-grounded refinement.
A two-lane flowchart depicts the data-prompt co-evolution loop. The upper lane represents Data (Test Inputs). It begins with an initial set of test items and adds new cases through a human-in-the-loop process. This results in a larger set that covers diverse, challenging, and ambiguous cases. The lower lane represents Prompt (Prompt Instructions). It starts with the current prompt and is refined through the same human-in-the-loop process. These revisions better reflect nuanced, application-specific policies. Centered between the lanes, the process is illustrated in which a machine synthesizes inputs, a human reviews model failures and articulates rationales, and the outcomes are routed in two directions: adding examples into the test set and revising the prompt.
@@ 1. Introduction
Large Language Models (LLMs) are driving a new wave of application development. This empowers people, including those without programming expertise, to shape application behavior through natural language instructions, a practice known as prompt engineering (Zamfirescu-Pereira et al., 2023; Subramonyam et al., 2025). However, it comes with an immediate challenge: a general-purpose LLM inherently lacks the local context and nuanced policies of a specific application. For instance, a small online community may wish to foster playful banter by allowing specific slang, a nuance lost on a strict, one-size-fits-all content moderation filter from a major tech vendor, such as OpenAI and Google. This gap forces people into a reactive cycle of prompt tinkering, making small, ad-hoc edits and spot-checking them against a few examples to patch failures at the margins.
@@
This ad-hoc approach reflects a deeper mismatch between the new speed of prompt iteration and the old tools for evaluation. Many current practices, such as building static test sets, are holdovers from the traditional machine learning era (Sambasivan et al., 2021). Back then, the slow, costly process of model fine-tuning necessitated a strict separation between data curation and model refinement. In today’s interactive landscape, where application developers can iterate on prompts in seconds, this inherited separation between test data and prompt refinement has become a major source of friction, preventing a truly fluid and responsive development process.
@@
This friction is compounded by a new bottleneck: the sheer difficulty of authoring comprehensive test data for the long tail of modern LLM-based applications. Manually anticipating the diverse ways users will interact with generative systems is a formidable task. To compensate, specialized and labor-intensive efforts have emerged. Red teaming, for example, uncovers critical vulnerabilities but often operates as a siloed, episodic exercise whose findings rarely integrate back into a persistent test suite (Feffer et al., 2024; Kiela et al., 2021; Quaye et al., 2024). Similarly, policy specifications or “constitutions” provide high-level guidance but tend to be static artifacts, difficult to update against a continuous stream of new edge cases (Bai et al., 2022; Petridis et al., 2024). Whether through ad-hoc tinkering, episodic testing, or static policies, a crucial feedback loop between discovering failures and refining behavior remains broken. We argue that the problem is the separation of test data and prompt instruction, where prompting is fast, but evaluation is slow.
To address this, we introduce data-prompt co-evolution, a new iterative workflow where test data and prompt instructions are no longer separate artifacts but interdependent components that evolve together. From a current prompt, the system generates new test cases that reveal its limitations, while each discovered failure informs how the prompt should be refined. This creates a continuous feedback loop: prompts grow more robust through exposure to challenging test data, while test sets expand to cover edge cases and ambiguities. Unlike traditional approaches that treat test sets as static benchmarks and prompt engineering as isolated tinkering, our approach recognizes that effective LLM application development requires these artifacts to shape each other iteratively.
@@
To operationalize this workflow, we present an interactive system that supports users to refine prompt instructions and create test sets together for generative applications. Users begin with an initial policy and prompt. The system proposes candidate user inputs, including those likely to be ambiguous under the given instruction, and runs the target model to reveal behavior. When the output is incorrect, the system engages the user in a dialog: they mark the example and articulate a short rationale. To test generalization to other examples, the system generates small neighborhoods of similar cases for quick labeling. These labeled slices both grow the test set and inform revisions to the prompt. Users are then directed to compare model responses before and after revision, with an LLM-as-judge (Zheng et al., 2023) seeded by the user’s own labels. Crucially, all labeled data persist, yielding a living, incrementally expanding test set.
@@
Our approach is grounded in formative interviews with five practitioners, which revealed a clear need for workflows that better integrate systematic, data-driven evaluation into the development loop. We then evaluated our system in a controlled study (N=16) against a baseline. Our workflow led participants to produce more detailed instructions with explicit boundaries and examples, build larger and richer test sets rapidly and systematically, induced more consistent changes in model behavior, and reported higher satisfaction compared to a baseline.
@@
  • A new workflow of data-prompt co-evolution, positioning test-set growth and specification refinement as incrementally co-evolving artifacts in LLM application development, informed by formative interviews with practitioners.
  • An interactive system that embodies this structured workflow that supports edge case discovery, rationale elicitation, neighborhood probing, and regression testing against a living test set.
  • Empirical insights from a controlled lab study, demonstrating that our workflow produces more specific prompt instructions, larger test sets, better alignment between participants’ intentions and model behavior, and higher user satisfaction than a baseline.
Our work contributes to HCI and human-AI interaction by advancing methods for democratizing LLM development and enabling more responsible, policy-aware use of AI. By shifting from prompt tinkering or episodic red teaming toward continuous, human-in-the-loop test-set growth, we show how interactive systems can help LLM application developers who are not necessarily AI experts make sense of subtle application-specific norms, codify them incrementally, and build more robust and accountable applications.
@@ 2.1. Human-Centered Prompt and Policy Design
Recent HCI research on LLMs and prompt engineering has framed prompts as design artifacts that encode requirements and governance choices rather than mere strings to be optimized. For example, ConstitutionMaker (Petridis et al., 2024) demonstrates how user critiques can be transformed into enforceable principles, while ROPE (Ma et al., 2025) highlights the role of prompts as requirement specifications. Policy Maps (Lam et al., 2025) uses interactive visualizations over a two-dimensional projection space to support the authoring of if-then policy rules over high-level concepts. PromptHive (Reza et al., 2025) illustrates how domain experts can co-create prompts. Together, these projects emphasize that prompts are sites where human values and policies are concretized.
Complementing these policy-oriented perspectives, other work has built interactive systems to help practitioners explore, compare, and refine prompts. PromptIDE (Strobelt et al., 2023) provides visual analytics support for prompt iteration, EvalLM (Kim et al., 2024) introduces workflows for comparing different prompt versions through structured evaluation, and ChainForge (Arawjo et al., 2024) enables rapid experimentation with LLM prompt chains. In addition, researchers have explored alternative familiar form factors, spreadsheets, for supporting prompt iteration in the context of data labeling and classification (He et al., 2025; Qian et al., 2025).
In parallel, the Natural Language Processing (NLP) community has largely pursued automatic prompt optimization, treating prompt design as a black-box search problem (Cheng et al., 2024). Recent techniques include black-box optimization for alignment, multi-branch optimization strategies (Yang et al., 2024), and embedding-based prompt search (Cheng et al., 2023). In addition, Yu et al. (Yu et al., 2025) uses synthetic data to automatically expose prompt weaknesses and revise prompts without human input. However, purely automated approaches often struggle to align with complex, implicit human intentions or safety policies. Our work argues for bringing the human into the loop to define and refine these governing policies.
@@ 2.2. Interactive Machine Learning and Labeling
Interactive machine learning (IML) has long shown that humans and models can co-adapt through iterative data curation (Amershi et al., 2014; Dudley and Kristensson, 2018). The Machine Teaching paradigm argues for exposing teaching levers, so non-experts can shape model behavior with structured workflows (Simard et al., 2017).
Recent mixed-initiative tools leverage LLMs for annotation. For example, PaTAT supports qualitative coding by synthesizing interpretable rules from a user’s codes (Gebreegziabher et al., 2023), and Wang et al. (Wang et al., 2024a) proposes a collaboration workflow where LLMs produce labels and explanations, a verifier assesses their quality, and human annotators re-annotate lower-quality ones.
@@
In parallel, the data-centric AI movement argues that systematic data design and curation, in addition to model tuning, drive performance, advocating practices for engineering data quality (Whang et al., 2023). Weak-supervision frameworks like Snorkel (Ratner et al., 2017) excel at throughput and coverage, but often miss opportunities for interactive elicitation.
@@ 2.3. Data Iteration and Test-Driven Development
A large body of HCI work argues that the importance of datasets in machine learning development. Chameleon (Hohman et al., 2020) shows how practitioners can attribute performance changes to data, comparing features, splits, and successive dataset versions through interactive visualizations. A series of interview studies further document the difficulty and iterative nature of data work in machine learning development in practice (Sambasivan et al., 2021; Holstein et al., 2019; Crisan and Fiore-Gartland, 2021; Qian et al., 2025).
Behavioral testing frameworks operationalize this stance by turning errors into executable tests. CheckList (Ribeiro et al., 2020) provides tooling to ideate perturbations and generate broad suites of checks, showing that teams uncover substantially more actionable bugs than with ad-hoc evaluation. There is also prior work that uses domain-specific languages to define precise error groups and run counterfactual rewrites (Wu et al., 2019, 2020). Zeno (Cabrera et al., 2023a) develops an interactive web UI to slice data, specify behaviors, and track results across tasks. However, these frameworks primarily operate on static datasets or require significant manual effort to formulate test cases.
Recent interactive workflows iteratively surface underexplored input patterns. ScatterShot (Wu et al., 2023) proposes a closely related idea in a different problem setting, slicing an existing unlabeled corpus to select diverse few-shot demonstrations for in-context learning. Our work overlaps in that both iteratively propose candidate examples for annotation. However, ScatterShot operates over a fixed corpus, whereas our work considers a challenging scenario where such a corpus may not exist and new examples must be generated to uncover failure modes. In addition, our workflow abstracts observed failures into natural-language policies that guide subsequent iterations. To address cases with limited unlabeled data, AdaTest (Ribeiro and Lundberg, 2022) and Farsight (Wang et al., 2024b) also utilize LLMs to synthesize cases for debugging and safety auditing, respectively. Our work similarly uses interactive data synthesis to explore undercovered regions but distinguishes itself by targeting policy refinement. Rather than stopping at test case generation (AdaTest) or harm awareness (Farsight), our workflow asks users to articulate rationales for failures. They serve as hypotheses that are translated into updated prompt instructions, coupling synthesized failure-grounded slices with iterative policy updates in a bidirectional co-evolution loop.
@@ 2.4. Bidirectional Human-AI Alignment and Criteria Drift
2.5. Human Roles in Red Teaming
Red teaming has emerged as a central strategy for stress-testing generative models. Automated pipelines exist for generating adversarial inputs (Perez et al., 2022), but researchers emphasize their limited creativity and coverage, making human input indispensable (Feffer et al., 2024). To scaffold human creativity at scale, the Adversarial Nibbler challenge (Quaye et al., 2024) recruited diverse communities to expose vulnerabilities in text-to-image generation, based on the Dynabench (Kiela et al., 2021) project. In addition, Yeh et al. (Yeh et al., 2025) and Deng et al. (Deng et al., 2025) leverage visual analytics to help red-teaming professionals explore the space of adversarial prompts.
@@ 3. Formative Study
To ground our work in real-world practices, we conducted a formative study to understand the challenges industry practitioners face when developing and refining LLM-based applications. We sought to answer the following questions: how do teams currently discover, diagnose, and address failures in their LLM systems, and how do these failures feed into their processes for improving their applications?
@@ 3.1. Participants and Procedure
We conducted 60-minute semi-structured interviews over video conferencing with five practitioners from various industries. [^1] Seven participants were recruited for interviews, but one withdrew at the end of their session and was excluded prior to analysis. Thematic analysis was initially conducted with six interviews, but excerpts from one participant were later excluded due to a potential conflict of interest. We verified that the four themes remained consistent when considering the five remaining interviews. The participants came from diverse professional roles, including data scientists, software engineers, product managers, and AI researchers with two to eleven years of experience in their fields. They were actively building and deploying a range of LLM applications, including natural-language-to-SQL query systems, call center dialogue summarization and classification, banking chatbots, and AI-assisted planning applications. Our interviews focused on their day-to-day workflows for development, evaluation, and iteration. The interviews were recorded, transcribed, and coded through thematic analysis. All interviews were conducted in Korean, and quotes presented in the next subsection were translated into English by the authors. Each participant received a gift card (KRW 60,000; approximately USD 45) as compensation. The study was approved by our institution’s IRB.
@@ 3.2.1. Development is a Process of Constant, Unending Iteration
This cycle of iteration—collecting new cases, updating prompts, and testing again—was described as never-ending. Below, we take a closer look at how practitioners actually conduct this iterative process.